Two guns situated on the top of a hill of height 10m, fire one shot each with the same speed 8.66 m/sec at some interval of time. One gun fires horizontally and other fires upwards at an angle of 60 degree with the horizontal. The shots collide in air at a point P. Find (i) the time interval between the firing, and (ii) the coordinates of the point P. Take origin of the coordinate system at the foot of the hill right below the muzzle and trajectories in x-y plane.

Grade:

1 Answers

Aman Bansal
12 years ago

Dear Siddharth,

For collision to take place, Equate the horizontal displacements of both the projectile taking Gun 1 being shot after T time interval.

Gun 1:

Sx=u*(t-T)

Gun 2:

Sx=u/2(t)

u*(t-T)=u/2*t

t=2T

similarly on equating the vertical displacement

to get (A) 1 second (B) 5root(3),5

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert